OpenStudy (anonymous):

How does (n-1) * (sqrt{n-1} + sqrt{n-1}) = 1/ n^(3/2)

OpenStudy (campbell_st):

so is the question \[n^{-1} \times (\sqrt{n^{-1}} + \sqrt{n^{-1}})\]

OpenStudy (anonymous):

yes

OpenStudy (alekos):

For me the LHS comes to 2/(n^3/2)

OpenStudy (campbell_st):

well other than missing a negative in the power \[n^{-1} \times 2n^{-\frac{1}{2}} = 2n^{-\frac{3}{2}}~~or~~\frac{2}{n^\frac{3}{2}}\]